# What You Trust Is Insecure: Demystifying How Developers (Mis)Use Trusted Execution Environments in Practice

### v2

@@ IV-B RQ2: What are the prevailing patterns in TEE development?
  • Cryptographic Primitives (68.2%): Core operations such as symmetric/asymmetric encryption, digital signatures, and key management, often reimplemented manually using OpenSSL [opensslOpenSSL] instead of vendor-provided SDKs.
@@
  • Secure Communication and I/O Handling (16.7%): Implementing secure channel endpoints like TLS [apostolopoulos1999transport], message integrity verification, or input/output filtering.
@@
Pattern 1: Modularity Integration. TEE usage in real-world systems is typically modular. Although recent research prototypes [li2023survey, zheng2021survey, paju2023sok] have made running entire applications inside TEEs more practical, developers seldom migrate full software stacks. Instead, they encapsulate only critical components such as cryptographic primitives that require strong confidentiality guarantees. This modular design reflects a pragmatic understanding of TEE limitations, including constrained memory and context switching overhead [gentilal2017trustzone, paju2023sok, ma2024bitdb], balancing protection requirements against performance costs. It also highlights a divergence between design intent and real-world practice. While TEEs were originally envisioned as general purpose trusted containers [mcgillion2015open, arnautov2016scone], developers often employ them as lightweight trust anchors [vuillermoz2022analysis] within larger system architectures to safeguard specific assets or operations. For example, a privacy preserving analytics service may perform data aggregation outside the TEE while confining attestation within it, which simplifies development but fragments the overall trust boundary, potentially introducing new attack surfaces as data flows between trusted and untrusted components.
@@
Pattern 2: Reimplementation and Bypassing of Standard TEE APIs. Another key finding is that many TEE projects reimplement cryptographic functionalities such as encryption and hashing algorithms instead of relying on official SDK-provided APIs. This trend is particularly evident in Intel SGX-based projects but is also observed in ARM TrustZone. Our quantitative analysis shows that in total 78 out of 241 projects (32.4%) manually implement cryptographic operations within enclaves or TAs, including 53 out of 149 SGX projects (35.6%) and 25 out of 92 TrustZone projects (27.1%), such as AsusWRT Merlin [flashrouters_asus_merlin], ppfl [mo2021ppfl] and mbedTLS [githubGitHubMbedTLSmbedtls].
@@
Manual reimplementation can arise from various motivations. Through manual inspection of code comments, commit messages, and internal discussions, we qualitatively identified two main drivers: (1) The limited functionality or unclear documentation of official SDKs. For example, the Intel SGX SDK exposes only a subset of cryptographic primitives [intel_sgx_sdk], while OP-TEE developers frequently report compatibility issues when integrating cryptographic APIs with modern toolchains [githubOPTEEoptee_os]. As a result, developers often import third-party cryptographic libraries such as mbedTLS [githubGitHubMbedTLSmbedtls] and OpenSSL [opensslOpenSSL], or build lightweight wrappers around untrusted counterparts, thereby compromising the original isolation guarantees of TEEs; (2) The desire for greater flexibility and control over cryptographic implementations. Developers may choose to customize algorithms or port existing implementations to meet specific security or performance requirements that official SDKs may not support without extensive modification. For instance, developers frequently port OpenSSL [opensslOpenSSL] into enclaves to leverage its rich feature set and optimizations, as observed in projects such as ppfl and darknetz [mo2020darknetz, mo2021ppfl].
@@
An illustrative snippet from ta_sec_key in AsusWRT Merlin [flashrouters_asus_merlin], a popular router firmware that secures private keys within TrustZone, is shown in Listing 2. The snippet manually and correctly performs an AES-128 key expansion routine [lee2022efficient], yet it exposes several weaknesses when executed inside a TEE. It relies on table lookups that are not constant time and may leak information through cache or timing side channels [van2018foreshadow], even within enclaves or TAs. Moreover, the function lacks proper memory hygiene, leaving temporary buffers and round keys uncleared in memory [shu2017study], and it omits secure random initialization, which is essential for maintaining cryptographic strength [egele2013empirical].
Overall, this pattern reveals a gap between the design intent of TEEs and their practical use in real-world development. Since TEEs are often viewed as black boxes that inherently ensure security, developers may underestimate the complexity of implementing secure cryptographic functionalities within them. When encountering difficulties with SDK APIs, they often choose to reimplement components themselves to achieve greater flexibility. This gap highlights the need to improve the usability of TEE SDKs through richer reference implementations, clearer documentation, and automated analysis tools that can detect or warn against unsafe reimplementations. In addition, supporting modular and pluggable cryptographic backends could allow developers to retain flexibility while preserving the core security guarantees of TEEs [li2023survey].
@@ 15 RoundKey[i * 4 + j] = RoundKey[(i - 4) * 4 + j] ^ temp[j];
Listing 2: Manual AES Key Expansion in ta_sec_key [flashrouters_asus_merlin]
@@
  • Modularized Multi-Model Deployment. Projects such as DarkneTZ isolate multiple model variants (e.g., TinyYOLO [khokhlov2020tiny], MobileNet [sinha2019thin]) into separate enclave modules or TAs, enabling dynamic selection or replacement without redeploying the entire enclave.
  • Simplified Single-Model Enclaves. Lightweight CNNs such as MiniVGG [pyimagesearchMiniVGGNetGoing] are embedded as single-purpose inference TAs by projects like ShadowNet [sun2023shadownet]. These models are typically statically linked, quantized, or reduced in depth to fit within the limited enclave memory space.
  • Compiler-Aided Optimization. Projects like Enigma [li2022enigma] apply ahead-of-time (AOT) simplification [thom2018survey] and runtime code generation [chen2018tvm] to produce enclave-optimized inference binaries during training.
@@
Although these strategies differ in implementation, they share a common goal of adapting AI workloads to the unique constraints of TEEs, such as limited memory (e.g., the 128 MB limit of Intel SGX 1.0 [intel_sgx_sdk]) and the performance overhead introduced by enclave transitions [paju2023sok]. By analyzing the source code and documentation of these projects, we observed that developers consciously select strategies based on their use cases and deployment environments, focusing on simplicity and optimization to ensure that AI models can be executed efficiently within TEEs without excessive overhead or complexity. This trend reflects a growing maturity in confidential AI development and suggests that TEE-based AI deployment is evolving from ad hoc experimentation toward more systematic engineering practices. It also motivates further research into specialized TEE support for AI workloads, such as reducing context-switch costs [li2023survey] and enabling larger memory capacity and secure model updates [yuan2024hypertheft].
@@ IV-C RQ3: What insecure practices and pitfalls are prevalent in real-world TEE projects?
While TEEs provide hardware-enforced isolation, their security guarantees can still be undermined by insecure coding practices and configuration mistakes. Prior research has identified several common pitfalls in TEE usage, such as missing input validation [ma2025diting], side-channel vulnerabilities [van2018foreshadow], and flawed key management [liu2022extending]. However, these analyses typically focus on theoretical vulnerabilities or isolated case studies, and the prevalence as well as diversity of insecure practices in real-world TEE applications remain largely unexplored.
Categorization Process. To bridge this gap, we systematically examined the source code of 241 TEE projects to uncover recurring insecure or discouraged practices that deviate from established secure coding guidelines. Rather than conducting exhaustive vulnerability assessments, our goal is to identify commonly recurring patterns, assess their potential impact, and derive actionable recommendations for improving TEE security in practice. Because no mature static or dynamic analysis tools currently exist for systematically auditing TEE code, existing methods such as Ma et al. [ma2025diting] can only detect a limited number of isolated cases and lack perfect accuracy. Therefore, our investigation relies primarily on manual code inspection, supported by the customized static analysis introduced in Section III. Specifically, we again used our static analysis scripts that apply keyword matching, prefix detection, and directory pattern recognition to identify (i) the presence of TEE-specific API calls like TEE_ and sgx_, and (ii) the location of source files within TEE-specific directories (e.g., ta/, enclave/, sgx/). Using the matched code context of these TEE-specific artifacts, we then manually analyzed all functions and tagged those that exhibited insecure practices.
To ensure both completeness (minimizing false negatives, i.e., missed insecure cases) and accuracy (reducing false positives, i.e., benign examples mistakenly flagged as insecure), we designed a three-stage labeling protocol. First, two authors independently inspected each repository by following Intel SGX and OP-TEE documentation, official SDK coding guidelines, and prior TEE vulnerability taxonomies such as Ma et al. [ma2025diting]. Second, the reviewers conducted calibration sessions on a pilot subset to refine the taxonomy and ensure consistent interpretation of insecure patterns. Third, a third author performed cross-validation, focusing on ambiguous cases to confirm that each identified instance represented a genuine security concern rather than benign code. This categorization process took approximately 100 hours in total, which was longer than the time required for the other two RQs.
@@
  • Missing Input Validation: Memory buffers, key lengths, and message structures are used without proper bounds checking, leading to potential buffer overflows or parsing errors [ma2025diting].
@@
Key Insights. The prevalence of these insecure patterns reveals a critical gap between the security guarantees envisioned by TEE designers and the realities of software engineering practice. Developers often underestimate the sensitivity of in-enclave operations or place excessive trust in the assumed inviolability of TEE boundaries. Through manual inspection of code and internal discussions, we qualitatively argue that this misconception is likely rooted in several factors: (1) a lack of formal training and awareness of TEE-specific secure coding practices; (2) limited security-focused documentation and examples in official SDKs; and (3) the inherent complexity of TEE development, which lacks mature tooling and automated analysis support compared to traditional software development, as also highlighted by prior studies [li2023survey, ma2025diting]. These findings collectively underscore the urgent need for better developer education, improved SDK usability, and dedicated tooling support to help developers identify and avoid common pitfalls when building TEE-based applications.
@@
Fixed Initialization Vector Usage. The first case concerns improper cryptographic parameter handling. AES-GCM encryption [lee2022efficient] requires a unique IV for every encryption to ensure semantic security. However, as shown below, the BI-SGX project initializes the IV as an all-zero array, violating this requirement. This design flaw likely stems from a misunderstanding of IV requirements or from developers reusing initialization templates for convenience. It allows identical plaintext blocks encrypted under the same key and IV to produce identical ciphertexts, enabling attackers to infer data patterns or recover plaintexts [bock2016nonce]. Critically, this vulnerability arises even if the encryption key remains securely sealed within the TEE, demonstrating that hardware isolation alone cannot compensate for weak cryptographic practice.
@@ V Discussion
Security Viewpoint: Practical Deviations from Ideal Models. Many projects diverge from the secure programming principles envisioned by TEEs. Common problems include logging sensitive data such as keys or authentication tokens using debug macros like printf, skipping attestation or verification steps, and hardcoding secrets in trusted code. We also find frequent misuse of cryptographic APIs, such as reusingIVrs and re-implementing primitives [egele2013empirical, nadi2016jumping]. Similar patterns have been reported in other security contexts like cryptographic libraries [fahl2012eve, acar2016you], but their impact is more severe in TEEs because enclave code operates under high trust assumptions. Once a vulnerability is introduced, it undermines the very isolation guarantees the hardware is meant to enforce [chen2018sgxpectre]. These deviations often stem from a combination of insufficient security awareness and practical barriers. Many developers rely on reused testing code or incomplete examples, and some face difficulties using complex attestation or cryptographic APIs, which leads to shortcuts that compromise security.
@@
Usability Viewpoint: Re-Implementation as a Signal of Friction. Frequent reimplementation of core SDK functions such as TEE_AllocateOperation() and custom handling of sessions or memory suggest that current TEE SDKs present substantial usability challenges. Developers often modify or bypass provided abstractions to match real-world workflows, revealing a gap between SDK design and actual development needs [wijayarathna2019using]. Such behaviors are not necessarily misuse but signals of friction. Developers adapt APIs when abstractions are too rigid, documentation lacks clarity, or interfaces remain overly hardware-centric. To make TEEs more approachable, SDKs should emphasize simplicity, clear examples, and safer defaults that guide developers toward correct usage rather than penalize experimentation.
@@
Design Viewpoint: Emerging Good Practices. Despite the issues discussed above, we also observe promising design patterns, particularly in AI-related projects. Developers increasingly partition workloads into modular components, deploy lightweight neural models that fit within enclave memory, and apply compiler-assisted transformations to generate optimized enclave binaries [mo2020darknetz, li2022enigma]. These design choices reflect a growing understanding of TEE constraints. Rather than treating TEEs solely as secure vaults for protecting keys, developers are integrating them as essential components of system architectures to execute more complex tasks securely. This shift demonstrates a maturation in TEE usage, where developers achieve a more effective balance between performance, usability, and security, and it also suggests rising pressure in practice to deliver secure systems. In this context, implementations that encapsulate entire applications within TEEs can be viewed as pragmatic solutions to avoid partitioning.
@@
Future Directions. Our findings point to several directions for improvement. First, SDK designers should enhance developer support through clearer documentation, simplified cryptographic interfaces, and secure code examples and project templates for common TEE development scenarios. Second, tool developers can build program analyzers [ma2025diting] to detect TEE-specific insecure API usage, and other recurring issues. Finally, researchers can design developer-friendly frameworks that make TEEs easier to use without sacrificing performance or security, thereby narrowing the gap between TEE’s theoretical guarantees and their practical deployment. In summary, the effectiveness of TEEs depends not only on hardware isolation but also on how developers understand and adopt them. Bridging the gap between idealized security assumptions and practical development remains crucial, and our study provides empirical evidence to support progress toward this goal.

### v3

@@ IV-B RQ2: What are the prevailing patterns in TEE development?
  • Cryptographic Primitives (68.2%): Core operations such as symmetric/asymmetric encryption, digital signatures, and key management, often reimplemented manually using OpenSSL [50] instead of vendor-provided SDKs.
@@
  • Secure Communication and I/O Handling (16.7%): Implementing secure channel endpoints like TLS [2], message integrity verification, or input/output filtering.
@@
Pattern 1: Modularity Integration. TEE usage in real-world systems is typically modular. Although recent research prototypes [35, 73, 51] have made running entire applications inside TEEs more practical, developers seldom migrate full software stacks. Instead, they encapsulate only critical components such as cryptographic primitives that require strong confidentiality guarantees. This modular design reflects a pragmatic understanding of TEE limitations, including constrained memory and context switching overhead [20, 51, 41], balancing protection requirements against performance costs. It also highlights a divergence between design intent and real-world practice. While TEEs were originally envisioned as general purpose trusted containers [43, 4], developers often employ them as lightweight trust anchors [64] within larger system architectures to safeguard specific assets or operations. For example, a privacy preserving analytics service may perform data aggregation outside the TEE while confining attestation within it, which simplifies development but fragments the overall trust boundary, potentially introducing new attack surfaces as data flows between trusted and untrusted components.
@@
Pattern 2: Reimplementation and Bypassing of Standard TEE APIs. Another key finding is that many TEE projects reimplement cryptographic functionalities such as encryption and hashing algorithms instead of relying on official SDK-provided APIs. This trend is particularly evident in Intel SGX-based projects but is also observed in ARM TrustZone. Our quantitative analysis shows that in total 78 out of 241 projects (32.4%) manually implement cryptographic operations within enclaves or TAs, including 53 out of 149 SGX projects (35.6%) and 25 out of 92 TrustZone projects (27.1%), such as AsusWRT Merlin [18], ppfl [45] and mbedTLS [42].
@@
Manual reimplementation can arise from various motivations. Through manual inspection of code comments, commit messages, and internal discussions, we qualitatively identified two main drivers: (1) The limited functionality or unclear documentation of official SDKs. For example, the Intel SGX SDK exposes only a subset of cryptographic primitives [25], while OP-TEE developers frequently report compatibility issues when integrating cryptographic APIs with modern toolchains [49]. As a result, developers often import third-party cryptographic libraries such as mbedTLS [42] and OpenSSL [50], or build lightweight wrappers around untrusted counterparts, thereby compromising the original isolation guarantees of TEEs; (2) The desire for greater flexibility and control over cryptographic implementations. Developers may choose to customize algorithms or port existing implementations to meet specific security or performance requirements that official SDKs may not support without extensive modification. For instance, developers frequently port OpenSSL [50] into enclaves to leverage its rich feature set and optimizations, as observed in projects such as ppfl and darknetz [46, 45].
@@
An illustrative snippet from ta_sec_key in AsusWRT Merlin [18], a popular router firmware that secures private keys within TrustZone, is shown in Listing 2. The snippet manually and correctly performs an AES-128 key expansion routine [32], yet it exposes several weaknesses when executed inside a TEE. It relies on table lookups that are not constant time and may leak information through cache or timing side channels [62], even within enclaves or TAs. Moreover, the function lacks proper memory hygiene, leaving temporary buffers and round keys uncleared in memory [55], and it omits secure random initialization, which is essential for maintaining cryptographic strength [16].
Overall, this pattern reveals a gap between the design intent of TEEs and their practical use in real-world development. Since TEEs are often viewed as black boxes that inherently ensure security, developers may underestimate the complexity of implementing secure cryptographic functionalities within them. When encountering difficulties with SDK APIs, they often choose to reimplement components themselves to achieve greater flexibility. This gap highlights the need to improve the usability of TEE SDKs through richer reference implementations, clearer documentation, and automated analysis tools that can detect or warn against unsafe reimplementations. In addition, supporting modular and pluggable cryptographic backends could allow developers to retain flexibility while preserving the core security guarantees of TEEs [35].
@@ 15 RoundKey[i * 4 + j] = RoundKey[(i - 4) * 4 + j] ^ temp[j];
Listing 2: Manual AES Key Expansion in ta_sec_key [18]
@@
  • Modularized Multi-Model Deployment. Projects such as DarkneTZ isolate multiple model variants (e.g., TinyYOLO [29], MobileNet [56]) into separate enclave modules or TAs, enabling dynamic selection or replacement without redeploying the entire enclave.
  • Simplified Single-Model Enclaves. Lightweight CNNs such as MiniVGG [53] are embedded as single-purpose inference TAs by projects like ShadowNet [58]. These models are typically statically linked, quantized, or reduced in depth to fit within the limited enclave memory space.
  • Compiler-Aided Optimization. Projects like Enigma [33] apply ahead-of-time (AOT) simplification [59] and runtime code generation [12] to produce enclave-optimized inference binaries during training.
@@
Although these strategies differ in implementation, they share a common goal of adapting AI workloads to the unique constraints of TEEs, such as limited memory (e.g., the 128 MB limit of Intel SGX 1.0 [25]) and the performance overhead introduced by enclave transitions [51]. By analyzing the source code and documentation of these projects, we observed that developers consciously select strategies based on their use cases and deployment environments, focusing on simplicity and optimization to ensure that AI models can be executed efficiently within TEEs without excessive overhead or complexity. This trend reflects a growing maturity in confidential AI development and suggests that TEE-based AI deployment is evolving from ad hoc experimentation toward more systematic engineering practices. It also motivates further research into specialized TEE support for AI workloads, such as reducing context-switch costs [35] and enabling larger memory capacity and secure model updates [70].
@@ IV-C RQ3: What insecure practices and pitfalls are prevalent in real-world TEE projects?
While TEEs provide hardware-enforced isolation, their security guarantees can still be undermined by insecure coding practices and configuration mistakes. Prior research has identified several common pitfalls in TEE usage, such as missing input validation [40], side-channel vulnerabilities [62], and flawed key management [36]. However, these analyses typically focus on theoretical vulnerabilities or isolated case studies, and the prevalence as well as diversity of insecure practices in real-world TEE applications remain largely unexplored.
Categorization Process. To bridge this gap, we systematically examined the source code of 241 TEE projects to uncover recurring insecure or discouraged practices that deviate from established secure coding guidelines. Rather than conducting exhaustive vulnerability assessments, our goal is to identify commonly recurring patterns, assess their potential impact, and derive actionable recommendations for improving TEE security in practice. Because no mature static or dynamic analysis tools currently exist for systematically auditing TEE code, existing methods such as Ma et al. [40] can only detect a limited number of isolated cases and lack perfect accuracy. Therefore, our investigation relies primarily on manual code inspection, supported by the customized static analysis introduced in Section III. Specifically, we again used our static analysis scripts that apply keyword matching, prefix detection, and directory pattern recognition to identify (i) the presence of TEE-specific API calls like TEE_ and sgx_, and (ii) the location of source files within TEE-specific directories (e.g., ta/, enclave/, sgx/). Using the matched code context of these TEE-specific artifacts, we then manually analyzed all functions and tagged those that exhibited insecure practices.
To ensure both completeness (minimizing false negatives, i.e., missed insecure cases) and accuracy (reducing false positives, i.e., benign examples mistakenly flagged as insecure), we designed a three-stage labeling protocol. First, two authors independently inspected each repository by following Intel SGX and OP-TEE documentation, official SDK coding guidelines, and prior TEE vulnerability taxonomies such as Ma et al. [40]. Second, the reviewers conducted calibration sessions on a pilot subset to refine the taxonomy and ensure consistent interpretation of insecure patterns. Third, a third author performed cross-validation, focusing on ambiguous cases to confirm that each identified instance represented a genuine security concern rather than benign code. This categorization process took approximately 100 hours in total, which was longer than the time required for the other two RQs.
@@
  • Missing Input Validation: Memory buffers, key lengths, and message structures are used without proper bounds checking, leading to potential buffer overflows or parsing errors [40].
@@
Key Insights. The prevalence of these insecure patterns reveals a critical gap between the security guarantees envisioned by TEE designers and the realities of software engineering practice. Developers often underestimate the sensitivity of in-enclave operations or place excessive trust in the assumed inviolability of TEE boundaries. Through manual inspection of code and internal discussions, we qualitatively argue that this misconception is likely rooted in several factors: (1) a lack of formal training and awareness of TEE-specific secure coding practices; (2) limited security-focused documentation and examples in official SDKs; and (3) the inherent complexity of TEE development, which lacks mature tooling and automated analysis support compared to traditional software development, as also highlighted by prior studies [35, 40]. These findings collectively underscore the urgent need for better developer education, improved SDK usability, and dedicated tooling support to help developers identify and avoid common pitfalls when building TEE-based applications.
@@
Fixed Initialization Vector Usage. The first case concerns improper cryptographic parameter handling. AES-GCM encryption [32] requires a unique IV for every encryption to ensure semantic security. However, as shown below, the BI-SGX project initializes the IV as an all-zero array, violating this requirement. This design flaw likely stems from a misunderstanding of IV requirements or from developers reusing initialization templates for convenience. It allows identical plaintext blocks encrypted under the same key and IV to produce identical ciphertexts, enabling attackers to infer data patterns or recover plaintexts [7]. Critically, this vulnerability arises even if the encryption key remains securely sealed within the TEE, demonstrating that hardware isolation alone cannot compensate for weak cryptographic practice.
@@ V Discussion
Security Viewpoint: Practical Deviations from Ideal Models. Many projects diverge from the secure programming principles envisioned by TEEs. Common problems include logging sensitive data such as keys or authentication tokens using debug macros like printf, skipping attestation or verification steps, and hardcoding secrets in trusted code. We also find frequent misuse of cryptographic APIs, such as reusingIVrs and re-implementing primitives [16, 48]. Similar patterns have been reported in other security contexts like cryptographic libraries [17, 1], but their impact is more severe in TEEs because enclave code operates under high trust assumptions. Once a vulnerability is introduced, it undermines the very isolation guarantees the hardware is meant to enforce [11]. These deviations often stem from a combination of insufficient security awareness and practical barriers. Many developers rely on reused testing code or incomplete examples, and some face difficulties using complex attestation or cryptographic APIs, which leads to shortcuts that compromise security.
@@
Usability Viewpoint: Re-Implementation as a Signal of Friction. Frequent reimplementation of core SDK functions such as TEE_AllocateOperation() and custom handling of sessions or memory suggest that current TEE SDKs present substantial usability challenges. Developers often modify or bypass provided abstractions to match real-world workflows, revealing a gap between SDK design and actual development needs [66]. Such behaviors are not necessarily misuse but signals of friction. Developers adapt APIs when abstractions are too rigid, documentation lacks clarity, or interfaces remain overly hardware-centric. To make TEEs more approachable, SDKs should emphasize simplicity, clear examples, and safer defaults that guide developers toward correct usage rather than penalize experimentation.
@@
Design Viewpoint: Emerging Good Practices. Despite the issues discussed above, we also observe promising design patterns, particularly in AI-related projects. Developers increasingly partition workloads into modular components, deploy lightweight neural models that fit within enclave memory, and apply compiler-assisted transformations to generate optimized enclave binaries [46, 33]. These design choices reflect a growing understanding of TEE constraints. Rather than treating TEEs solely as secure vaults for protecting keys, developers are integrating them as essential components of system architectures to execute more complex tasks securely. This shift demonstrates a maturation in TEE usage, where developers achieve a more effective balance between performance, usability, and security, and it also suggests rising pressure in practice to deliver secure systems. In this context, implementations that encapsulate entire applications within TEEs can be viewed as pragmatic solutions to avoid partitioning.
@@
Future Directions. Our findings point to several directions for improvement. First, SDK designers should enhance developer support through clearer documentation, simplified cryptographic interfaces, and secure code examples and project templates for common TEE development scenarios. Second, tool developers can build program analyzers [40] to detect TEE-specific insecure API usage, and other recurring issues. Finally, researchers can design developer-friendly frameworks that make TEEs easier to use without sacrificing performance or security, thereby narrowing the gap between TEE’s theoretical guarantees and their practical deployment. In summary, the effectiveness of TEEs depends not only on hardware isolation but also on how developers understand and adopt them. Bridging the gap between idealized security assumptions and practical development remains crucial, and our study provides empirical evidence to support progress toward this goal.